OpenStudy (anonymous):

Enter the value of x such that ³√3ˣ multiplied by 3 and 5/3 equals ³√3⁷.

OpenStudy (isaiah.feynman):

You mean?\[\sqrt[3]{3^{x}}\times3\frac{ 5 }{ 3 } = \sqrt[3]{3^{7}}\]

OpenStudy (anonymous):

Yes!

jhonyy9 (jhonyy9):

so you know that 5 3--- = 5 3 after this how you can continue it ?

jhonyy9 (jhonyy9):

cuberoot(3^x) = 3^(x/3) on the other side there is cuberoot(3^7) = 3^(7/3) so than you can rewriting your equation in this form : 3^(x/3) *5 = 3^(7/3) do you can continue from this ?